OpenStudy (anonymous):

@kirbykirby The graph plots four equations, A, B, C, and D: Line A joins ordered pair negative 6, 16 and 9, negative 4. Line B joins ordered pair negative 2, 20 and 8, 0. Line C joins ordered pair negative 7, negative 6 and 6, 20. Line D joins ordered pair 7, 20 and 0, negative 7. Which pair of equations has (4, 8) as its solution? Equation A and Equation C Equation B and Equation C Equation C and Equation D Equation B and Equation D

OpenStudy (anonymous):

@phi @inkyvoyd Please Help!

OpenStudy (kirbykirby):

find the equation of the lines : that is, you can find the line in the form y=mx+b easily by finding the slope "m" first... the slope formula being \[\frac{y_2-y_1}{x_2-x_1}\] where the numerator values are the y-coordinates of your points, and the denominator are the x-coordinates of your points. Once you have "m" you just need to find the y-intercept "b" which you can do by plugging in either of your 2 points into the formula y = mx+b and then solve for "b" You should do this for all the lines A, B, C, D Finally, Just plug in the values of x and y into the line equation for the point (4, 8). If the equation is satisfied (i.e. you get something like "8 = 8") then the point will be on the line . You'll have to try plugging in (4, 8) into all equations and see if the point satisfies the equation

OpenStudy (anonymous):

For A I got -20 over 3. For B, -2. For C, 2. And for D, 27 over 7

OpenStudy (anonymous):

That are the slopes. I am confused on what to do next?

OpenStudy (anonymous):

Or well, actually HOW to do it. I know I need to find the Y-intercept but I am having difficulties doing so.

OpenStudy (kirbykirby):

Once you have the slopes, say for A you got -20/3, then you have this as your equation so far: \[y = \frac{-20}{3}x+b \] So you can pick one point (either one of the two is okay), and plug the x and y-coordinates where you see "x" and "y" into the equation, this will allow you to find the y-intercept "b"

OpenStudy (anonymous):

So, \[16=\frac{ -20 }{ 3 }(6)+b\]

OpenStudy (kirbykirby):

that 6 should be " - 6"

OpenStudy (anonymous):

Oh okay. So instead it would be\[16=\frac{ -20 }{ 3 }(-6)+b\]

OpenStudy (anonymous):

But now that I have the equation for A, what is my next step. How do I find the value of B?

OpenStudy (anonymous):

Not line B but the (+b) in the equation

OpenStudy (kirbykirby):

You can solve for it algebraically.. \[16 = \frac{-20}{3}(-6)+b\\ 16 = \frac{120}{3}+b \]... can you find b? (find the value of 120/3 = ?, and subtract both sides by that value and you will have isolated "b")

OpenStudy (anonymous):

-24?

OpenStudy (kirbykirby):

yup!

OpenStudy (anonymous):

And B would be:\[20=-2(-2)+b\]

OpenStudy (kirbykirby):

yes

OpenStudy (anonymous):

\[16=4+b\]

OpenStudy (anonymous):

b=12?

OpenStudy (kirbykirby):

The last step you just have to verify if (4,8) satisfies the equations, so like 8 = m(4) + b so you need to check if your right side of the equation will equal 8 :)

OpenStudy (kirbykirby):

yes

OpenStudy (kirbykirby):

I hope this gives you enough to finish the exercise as I have to go :o!

OpenStudy (anonymous):

It does! Thank You!

OpenStudy (kirbykirby):

yw :) ! byee